OpenStudy (anonymous):

-9x _ 5= -95 A) 17 B)11 C)10 D) -10

OpenStudy (haseeb96):

-9x-5=-95 -9x=-95+5 -9x=-90 multiply both sides by -1 9x=90 x=90\9 x=10 this is the correct answer

OpenStudy (anonymous):

Thanks! Can you help me with another one?